# Improving Steering Vectors by Targeting Sparse Autoencoder Features

### v1

@@ Abstract
To control the behavior of language models, steering methods attempt to ensure that outputs of the model satisfy specific pre-defined properties. Adding steering vectors to the model is a promising method of model control that is easier than finetuning, and may be more robust than prompting.However, it can be difficult to anticipate the effects of steering vectors produced by almost all existing methods, such as CAA (Panickssery et al., 2024) or the direct use of SAE latents (Templeton et al., 2024). In our work, we address this issue by using SAEs to measure the effects of steering vectors, giving us a method that can be used to understand the causal effect of any steering vector intervention. We use this method for measuring causal effects to develop an improved steering method, SAE-Targeted Steering (SAE-TS), which finds steering vectors to target specific SAE features while minimizing unintended side effects. We show that overall, SAE-TS balances steering effects with coherence better than CAA and SAE feature steering, when evaluated on a range of tasks.
@@ 2 Related work
Mechanistic Interpretability and SAEs. Mechanistic Intepretability aims to understand how LLMs function internally by breaking down the models into understandable components (Olah et al., 2020; Olah, ). The recent development of sparse autoencoders (Ng, 2011; Bricken et al., 2023; Cunningham et al., 2023) gives us a way to decompose the residual stream of transformer models into sparse and often human understandable features. Sparse autoencoders provide evidence supporting the hypothesis that the latent space of LLMs is composed of linear and interpretable directions (Arora et al., 2018; Elhage et al., 2022). Templeton et al. (2024) demonstrated that these directions found by the SAEs can be used for steering.
@@ 5 Evaluations
Steering is sensitive to the norm of the steering vector so we vary the scale $\alpha$ and plot the three scores for our SAE-TS steering method as shown in Figure 2. Starting from scale $\alpha=0$ (the unsteered model), the Coherence score is 0.64 while the Behavioral score is 0, making the Behavioral*Coherence score 0. As we increase the scale, the Behavioral score either keeps rising to more than 0.9 in the case of the "London" task, or peaks at around 0.77 and then goes back down as is the case in the "Wedding" task. Interestingly, the Coherence score first rises slightly to a peak of around 0.71 and then drops off. This rising Coherence score phenomenon is specific to the SAE-TS method and we explore it further in Section A.3.
@@
Figure 3 shows the performance of these methods across the 9 steering tasks. The $x$-axis represents the scaling factor $\alpha$ while the $y$-axis is Behavioral*Coherence score. Looking at the peak of the score (how good is it when the appropriate scale is chosen), displayed in Table 2, we see that the our SAE-TS method outperforms the baseline methods on 7 of the 9 tasks. The two exceptions are the "French" task, where CAA is the best method, and the "praise" task, where SAE steering is about as good as SAE-TS.
@@
To verify our findings across model sizes, we also conducted the same analysis using the Gemma-2-9B model and present these results in Appendix C. We observe broadly similar trends, with SAE-TS outperforming the baseline methods across most tasks.
@@
Anger 0.0976 0.0732 0.2302
Christian 0.0346 0.0901 0.3445
Conspiracy 0.1150 0.2097 0.3858
French 0.3324 0.0586 0.3040
London 0.0093 0.0073 0.5753
Love 0.1427 0.1082 0.4431
Praise 0.1291 0.2788 0.2754
Want to die 0.0323 0.0872 0.1992
Wedding 0.2119 0.2389 0.5500
@@ Acknowledgements
This research was conducted as part of the MATS 6.0 program. We would like to thank Yeu-Tong Lau and Kai Fronsdal for valuable discussions that helped develop the SAE-TS method.
@@ A.1 Empirical justification
Anger 0.2302 0.1790
Christian 0.3445 0.0008
Conspiracy 0.3858 0.0468
French 0.3040 0.1966
London 0.5753 0.0504
Love 0.4431 0.2473
Praise 0.2754 0.0679
Want to die 0.1992 0.1630
Wedding 0.5500 0.0492
@@ A.3 Effect of the bias term
We suspect that the the initial Coherence bump that we observe in Section 5 is due to the bias term ($-\bm{Mb}$) in Equation 3. We can see this in Figure 5 where we plot the Coherence score when steering with just $-\bm{Mb}$ against steering with random normal vectors.
@@ Appendix C Gemma-2-9B Results
We perform the same evaluations as in Section 5 for the gemma-2-9B model and plot the results in Figure 6. We use the Gemmascope layer 12 residual stream SAE with width 16k and an L0 of 130. It was easy to find relevant features for all the tasks except for the London task. There is no feature in this particular SAE which fires only on London, thus we use a mix of the following three features: 6915 - large cities including London, 9983 - cities, 11748 - United Kingdom. We see that SAE-TS is once again the best steering method achieving higher Behavioral*Coherence scores than either CAA or SAE steering in 8 out of the 9 tasks.
@@ Appendix D Rotation Steering
Anger 0.0976 0.0732 0.2302 0.1790 0.2380
Christian 0.0346 0.0901 0.3445 0.0008 0.1022
Conspiracy 0.1150 0.2097 0.3858 0.0468 0.4008
French 0.3324 0.0586 0.3040 0.1966 0.3717
London 0.0093 0.0073 0.5753 0.0504 0.2466
Love 0.1427 0.1082 0.4431 0.2473 0.3618
Praise 0.1291 0.2788 0.2754 0.0679 0.2938
Want to die 0.0323 0.0872 0.1992 0.1630 0.2277
Wedding 0.2119 0.2389 0.5500 0.0492 0.5305
@@ I.2 CAA London
Scale = 160. It does not mention London directly, but does talk about the Thames.
"<bos>I think the people on the other end of the phone is for support but not for our
@@
"<bos>I think it is nice but we have only just visited once, we have never been on
    the cruise this, you have to be a real ale one I can",
"<bos>I think in a few short months it will be better than the Thames boat trip.
@@
"<bos>I think it would be cool if you could take a bus that takes you to the wharf
    for a pint on a Sunday morning and take a walk to the harbour",

### v2

@@ Abstract
To control the behavior of language models, steering methods attempt to ensure that outputs of the model satisfy specific pre-defined properties. Adding steering vectors to the model is a promising method of model control that is easier than finetuning, and may be more robust than prompting.However, it can be difficult to anticipate the effects of steering vectors produced by methods such as CAA (Panickssery et al., 2024) or the direct use of SAE latents (Templeton et al., 2024). In our work, we address this issue by using SAEs to measure the effects of steering vectors, giving us a method that can be used to understand the causal effect of any steering vector intervention. We use this method for measuring causal effects to develop an improved steering method, SAE-Targeted Steering (SAE-TS), which finds steering vectors to target specific SAE features while minimizing unintended side effects. We show that overall, SAE-TS balances steering effects with coherence better than CAA and SAE feature steering, when evaluated on a range of tasks.
@@ 2 Related work
Mechanistic Interpretability and SAEs. Mechanistic Intepretability aims to understand how LLMs function internally by breaking down the models into understandable components (Olah et al., 2020; Olah, ). The recent development of sparse autoencoders (Ng, 2011; Bricken et al., 2023; Cunningham et al., 2023) gives us a way to decompose the residual stream of transformer models into sparse and often human understandable features. Sparse autoencoders provide evidence supporting the hypothesis that the latent space of LLMs is composed of linear and interpretable directions (Arora et al., 2018; Elhage et al., 2022). Templeton et al. (2024) and Zhao et al. (2024) demonstrate that these directions found by the SAEs can be used for steering.
@@ 5 Evaluations
Steering is sensitive to the norm of the steering vector so we vary the scale $\alpha$ and plot the three scores for our SAE-TS steering method as shown in Figure 2. Starting from scale $\alpha=0$ (the unsteered model), the Coherence score is 0.64 while the Behavioral score is 0, making the Behavioral*Coherence score 0. As we increase the scale, the Behavioral score either keeps rising to more than 0.9 in the case of the "London" task, or peaks at around 0.75 and then goes back down as is the case in the "Wedding" task. Interestingly, the Coherence score first rises slightly to a peak of around 0.71 and then drops off. This rising Coherence score phenomenon is specific to the SAE-TS method and we explore it further in Section A.3.
@@
Figure 3 shows the performance of these methods across the 9 steering tasks. The $x$-axis represents the scaling factor $\alpha$ while the $y$-axis is Behavioral*Coherence score. Looking at the peak of the score (how good is it when the appropriate scale is chosen), displayed in Table 2, we see that the our SAE-TS method outperforms the baseline methods on 7 of the 9 tasks. The two exceptions are the "Christian" task, where CAA is the best method, and the "Conspiracy" task, where CAA is tied with SAE-TS.
@@
While we primarily focus on Gemma-2-2B in this paper, we also validate our methods on Gemma-2-9B and present these results in Appendix C. We observe broadly similar trends, with SAE-TS outperforming the baseline methods on average, though we see that CAA sometimes performs better on the larger model.
@@
Anger 0.1666 0.0810 0.2424
Christian 0.3486 0.0902 0.3302
Conspiracy 0.3748 0.2319 0.3729
French 0.2684 0.0589 0.3186
London 0.0476 0.0061 0.5380
Love 0.3039 0.1004 0.4234
Praise 0.1332 0.2654 0.2842
Want to die 0.1284 0.0649 0.1870
Wedding 0.1768 0.2626 0.5432
Average 0.2165 0.1290 0.3600
@@ Acknowledgements
This research was conducted as part of the MATS 6.0 program. We would like to thank Yeu-Tong Lau and Kai Fronsdal for valuable discussions that helped develop the SAE-TS method. We would also like to thank the GitHub user https://github.com/spfaul for finding a bug in our CAA implementation which we have fixed and regenerated results for, in this version of the paper.
@@ A.1 Empirical justification
Anger 0.2424 0.1899
Christian 0.3302 0.0025
Conspiracy 0.3729 0.0644
French 0.3186 0.1912
London 0.5380 0.0722
Love 0.4234 0.2682
Praise 0.2842 0.0638
Want to die 0.1870 0.1683
Wedding 0.5432 0.0567
@@ A.3 Effect of the bias term
We suspect that the initial Coherence bump that we observe in Section 5 is due to the bias term ($-\bm{Mb}$) in Equation 3. We can see this in Figure 5 where we plot the Coherence score when steering with just $-\bm{Mb}$ against steering with random normal vectors.
@@ Appendix C Gemma-2-9B Results
We perform the same evaluations as in Section 5 for the gemma-2-9B model and plot the results in Figure 6 and Table 5. We use the Gemmascope layer 12 residual stream SAE with width 16k and an L0 of 130. It was easy to find relevant features for all the tasks except for the London task. There is no feature in this particular SAE which fires only on London, thus we use a mix of the following three features: 6915 - large cities including London, 9983 - cities, 11748 - United Kingdom. We see that SAE-TS achieves higher average Behavioral*Coherence score than either CAA or SAE steering. However, CAA beats SAE-TS in 4 of the 9 tasks and ties in one of them.
@@
Steering Goal CAA SAE SAE-TS (ours)
Anger 0.2336 0.1317 0.2310
Christian 0.3162 0.1548 0.2769
Conspiracy 0.4343 0.3003 0.3315
French 0.3203 0.3272 0.4196
London 0.0656 0.1118 0.3135
Love 0.3373 0.1160 0.2710
Praise 0.1965 0.3743 0.5188
Want to die 0.1521 0.1646 0.1265
Wedding 0.2216 0.2851 0.5494
Average 0.2531 0.2184 0.3376
Table 5: Gemma-2-9B maximum Behavioral*Coherence scores.
@@ Appendix D Rotation Steering
Anger 0.1666 0.0810 0.2424 0.1899 0.2460
Christian 0.3486 0.0902 0.3302 0.0025 0.0870
Conspiracy 0.3748 0.2319 0.3729 0.0644 0.4074
French 0.2684 0.0589 0.3186 0.1912 0.3724
London 0.0476 0.0061 0.5380 0.0722 0.2745
Love 0.3039 0.1004 0.4234 0.2682 0.3851
Praise 0.1332 0.2654 0.2842 0.0638 0.2833
Want to die 0.1284 0.0649 0.1870 0.1683 0.2167
Wedding 0.1768 0.2626 0.5432 0.0567 0.5313
Average 0.2165 0.1290 0.3600 0.1197 0.3115
@@ I.2 CAA London
Scale = 100.
<bos>I think they are a bit misleading as the term "London" as a hotel may be in poor
@@
<bos>I think the biggest challenge this year for the FA has been what they do with the
    2nd stage teams. The clubs have been locked away from each other,
<bos>I think I read somewhere that they’ve been asked to be more involved in the music
    scene before. I mean I get that the band probably won’t,
<bos>I think that people should feel ’theatred’ rather than ’thespian’ and ’theatre’